OpenStudy (anonymous):

Show graphically that the system of linear equations 2x+4y=10 and 3x + 6y = 12 has no solution. Explain why.

OpenStudy (anonymous):

2x + 4y = 10 * 3 = 6x + 12y = 30 3x + 6y = 12 * 2 = 6x + 12y = 24 as the two variables with the same coefficients have two different values, there is no possible solution to the system of linear equations

OpenStudy (anonymous):

graphically , the lines are parallel and dont no not intersect , thats why there is no solution to the simultaneous equations